OpenStudy (anonymous):

Solve: 25x2 − 36 = 0. x = 6, 5 x = −6, −5 x = 36 over 25, x = − 36 over 25 x = 6 over 5, x = − 6 over 5

OpenStudy (anonymous):

@Luigi0210

Nnesha (nnesha):

difference of squares \[\huge\rm a^2-b^2=(a+b)(a-b)\]

Nnesha (nnesha):

take square root of both terms (sqrt of 1st term + sqrt of 2nd term)(sqrt of 1st term - sqrt of 2nd term)

OpenStudy (anonymous):

I am lost

Nnesha (nnesha):

here is an example if i have 4 i can write it as 2^2 power right ? so what number to the 2 power = 25

OpenStudy (anonymous):

5

Nnesha (nnesha):

yes right what about 36 ?

OpenStudy (anonymous):

6

Nnesha (nnesha):

right so you can write it as \[\huge\rm 5^2x^2-6^2\] which is same as \[\rm a^2-b^2=(a-b)(a+b)\]

Nnesha (nnesha):

a is first term and b is 2nd term what is a and b in ur equation ?

OpenStudy (anonymous):

a=5^2 and b=6^2

Nnesha (nnesha):

not just 5^2 a=5^2x^2 so write it as (a-b)(a+B)

Nnesha (nnesha):

ohh wait

Nnesha (nnesha):

a= 5x b=6 not squares

Nnesha (nnesha):

difference of squares \[\huge\rm\color{reD}{ a}^2-\color{blue}{b}^2=(\color{Red}{a}+\color{blue}{b})(\color{reD}{a}-\color{blue}{b})\]

Nnesha (nnesha):

well guess what its easy to take square roots